OpenStudy (anonymous):

Find the slope of the line that passes through the points (5, 8) and (9, 4). A. 3 B. 1 C. -1 D. four-thirds

OpenStudy (texaschic101):

you will have to use the slope formula slope = (y2 - y1) / (x2 - x1) (5,8) x1 = 5 and y1 = 8 (9,4) x2 = 9 and y2 = 4 now we sub slope = (4 - 8) / (9 - 5) slope = -4/4 slope = -1 any questions ?